# A Case Study on Remediating Math Mistakes

### v2

@@ 1 Introduction
[…]
@@ 6 Evaluation
lets_closer 2.42 steps_took 1.93 explain_arrived 3.71 start_remember 1.73
closer_look 2.42 thought_process 1.54 can_explain 3.59 good_try 1.69
appreciate_effort 2.23 review_concept 1.54 arrived_answer 2.99 thats_good 1.59
effort_lets 2.23 took_arrive 1.43 are_sure 2.24 thats_right 1.59
look_problem 2.01 explain_steps 1.43 sure_that 2.24 try_again 1.39
correct_solution 1.86 help_understand 1.30 arrived_number 1.73 good_start 1.37
sure_weve 1.61 understand_concept 1.30 and_long 1.41 lets_break 1.24
good_attempt 1.61 number_sides 1.30 answer_step 1.41 example_apples 1.22
lets_steps 1.61 lets_step 1.26 numbers_sure 1.41 represents_glasses 1.22
great_effort 1.33 explain_thought 1.16 can_tell 1.33 original_number 1.22
@@ Acknowledgements
We’d like to thank Jiang Wu, Hannah Shuchhardt, and two anonymous individuals for their help and feedback on our work; the Stanford NLP group, Joy He-Yueya, Gabriel Poesia, Myra Cheng, Kristina Gligorić, Ali Malik and Roma Patel for their feedback on the paper; Jesse Mu for pointing us to Sankey diagrams; and IMK for the Bridge inspiration.
@@ D.1 No Decision-Making Condition
No Decision-Making Prompt for gpt-4 and gpt-3.5-turbo You are an experienced elementary math teacher and you are going to respond to a student’s mistake in a useful and caring way. The problem your student is solving is on topic: {lesson_topic}.
{c_h}
tutor (maximum one sentence):
@@
No Decision-Making Prompt for llama-2 ### System:
You are an experienced elementary math teacher and you are going to respond to a student’s mistake in a useful and caring way.
### User:
Lesson topic: {lesson_topic}
Conversation:
{c_h}
### Assistant:
tutor (maximum one sentence):
@@ D.2 Expert Decision-Making Condition
Decision-Making Prompt for gpt-4 and gpt-3.5-turbo You are an experienced elementary math teacher and you are going to respond to a student’s mistake in a useful and caring way. The problem your student is solving is on topic: {lesson_topic}. {e} {z_what} in order to {z_why}.
{c_h}
tutor (maximum one sentence):
@@
Decision-Making Prompt for llama-2 ### System:
You are an experienced elementary math teacher and you are going to respond to a student’s mistake in a useful and caring way.
### User:
{e} {z_what} in order to {z_why}.
Lesson topic: {lesson_topic}
Conversation:
{c_h}
### Assistant:
tutor (maximum one sentence):
@@ D.3 Self decision-making condition
Determine Error ($e$) with gpt-4 and gpt-3.5-turbo. You are an experienced elementary math teacher. Your task is to read a conversation snippet of a tutoring session between a student and tutor, and determine what type of error the student makes in the conversation. We have a list of common errors that students make in math, which you can pick from. We also give you the option to write in your own error type if none of the options apply.
Error list:
0. Student does not seem to understand or guessed the answer.
1. Student misinterpreted the question.
2. Student made a careless mistake.
3. Student has the right idea, but is not quite there.
4. Student’s answer is not precise enough or the tutor is being too picky about the form of the student’s answer.
5. None of the above, but I have a different description (please specify in your reasoning).
6. Not sure, but I’m going to try to diagnose the student.
Here is the conversation snippet:
Lesson topic: {lesson_topic}
Conversation:
{c_h}
Why do you think the student made this mistake? Pick an option number from the error list and provide the reason behind your choice. Format your answer as: [{"answer": #, "reason": "write out your reason for picking # here"}]
@@
Determine Error ($e$) with llama-2. ### System:
You are an experienced elementary math teacher. Your task is to read a conversation snippet of a tutoring session between a student and tutor, and determine what type of error the student makes in the conversation. We have a list of common errors that students make in math, which you can pick from. We also give you the option to write in your own error type if none of the options apply.
Error list:
0. Student does not seem to understand or guessed the answer.
1. Student misinterpreted the question.
2. Student made a careless mistake.
3. Student has the right idea, but is not quite there.
4. Student’s answer is not precise enough or the tutor is being too picky about the form of the student’s answer.
5. None of the above, but I have a different description (please specify in your reasoning).
6. Not sure, but I’m going to try to diagnose the student.
Format your answer as: [{"answer": #, "reason": "write out your reason for picking # here"}]
### User:
Lesson topic: {lesson_topic}
Conversation:
{c_h}
### Assistant:
[{"answer":
@@
Determine Strategy and Intention ($z_{\text{what}},z_{\text{why}}$) with gpt-4 and gpt-3.5-turbo. You are an experienced elementary math teacher. Your task is to read a conversation snippet of a tutoring session between a student and tutor where a student makes a mistake. You should then determine what strategy you want to use to remediate the student’s error, and state your intention in using that strategy. We have a list of common strategies and intentions that teachers use, which you can pick from. We also give you the option to write in your own strategy or intention if none of the options apply.
Strategies:
0. Explain a concept
1. Ask a question
2. Provide a hint
3. Provide a strategy
4. Provide a worked example
5. Provide a minor correction
6. Provide a similar problem
7. Simplify the question
8. Affirm the correct answer
9. Encourage the student
10. Other (please specify in your reasoning)
Intentions:
0. Motivate the student
1. Get the student to elaborate their answer
2. Correct the student’s mistake
3. Hint at the student’s mistake
4. Clarify a student’s misunderstanding
5. Help the student understand the lesson topic or solution strategy
6. Diagnose the student’s mistake
7. Support the student in their thinking or problem-solving
8. Explain the student’s mistake (eg. what is wrong in their answer or why is it incorrect)
9. Signal to the student that they have solved or not solved the problem
10. Other (please specify in your reasoning)
Here is the conversation snippet:
Lesson topic: {lesson_topic}
Conversation:
{c_h}
How would you remediate the student’s error and why? Pick the option number from the list of strategies and intentions and provide the reason behind your choices. Format your answer as: [{"strategy": #, "intention": #, "reason": "write out your reason for picking that strategy and intention"}]
@@ 10. Other (please specify in your reasoning)
Determine Strategy and Intention ($z_{\text{what}},z_{\text{why}}$) with llama-2. ### System:
You are an experienced elementary math teacher. Your task is to read a conversation snippet of a tutoring session between a student and tutor where a student makes a mistake. You should then determine what strategy you want to use to remediate the student’s error, and state your intention in using that strategy. We have a list of common strategies and intentions that teachers use, which you can pick from. We also give you the option to write in your own strategy or intention if none of the options apply.
Strategies:
0. Explain a concept
1. Ask a question
2. Provide a hint
3. Provide a strategy
4. Provide a worked example
5. Provide a minor correction
6. Provide a similar problem
7. Simplify the question
8. Affirm the correct answer
9. Encourage the student
10. Other (please specify in your reasoning)
Intentions:
0. Motivate the student
1. Get the student to elaborate their answer
2. Correct the student’s mistake
3. Hint at the student’s mistake
4. Clarify a student’s misunderstanding
5. Help the student understand the lesson topic or solution strategy
6. Diagnose the student’s mistake
7. Support the student in their thinking or problem-solving
8. Explain the student’s mistake (eg. what is wrong in their answer or why is it incorrect)
9. Signal to the student that they have solved or not solved the problem
10. Other (please specify in your reasoning)
Format your answer as: [{"strategy": #, "intention": #, "reason": "write out your reason for picking # here"}]
### User:
Lesson topic: {lesson_topic}
Conversation:
{c_h}
### Assistant:
[{"strategy":
@@ Appendix F Lexical analysis
effort_remember 2.30 can_explain 2.01 actually_correct 2.73 thats_close 1.69
effort_double 1.36 great_start 1.88 correct_answer 1.93 example_help 1.58
clarify_mean 1.36 can_tell 1.85 number_baseballs 1.42 can_think 1.47
noticed_mistake 1.36 got_answer 1.63 problem_remember 1.42 can_try 1.47
great_effort 1.34 explain_got 1.53 job_attempting 1.13 good_start 1.36
@@
user_lesson 6.69 lets_closer 4.02 user_student 5.59 lets_closer 3.52
user_tutor 4.79 closer_look 4.02 student_responds 4.39 closer_look 3.52
teacher_going 3.03 look_problem 2.85 student_understand 3.67 right_track 2.79
experienced_elementary 3.03 problem_break 1.67 response_provide 3.52 youre_right 2.29
going_respond 3.03 groups_objects 1.54 help_student 3.14 student_mistake 1.93

### v3

@@ 1 Introduction
Figure 1: \⃝raisebox{-0.8pt}{1} Closing the knowledge gap at scale. LLMs and novice tutors lack the pedagogical knowledge to engage with student mistakes, yet they are readily available for 1:1 tutoring. Experts like experienced teachers have the pedagogical knowledge, but are hard to scale. \⃝raisebox{-0.8pt}{2} How do we model the expert’s thought process? Our work builds Bridge which leverage cognitive task analysis to translate the latent thought process of experts into a decision-making model. \⃝raisebox{-0.8pt}{3} Applying Bridge with LLMs. To bridge the knowledge gap, we scale the expert’s knowledge with LLMs using the expert-guided decision-making model.
@@ 6 Evaluation
lets_closer 2.76 steps_took 2.04 can_explain 4.98 good_try 1.82
closer_look 2.68 review_concept 1.66 explain_arrived 4.78 start_remember 1.58
effort_lets 2.55 understand_concept 1.56 arrived_answer 4.2 thats_right 1.57
appreciate_effort 2.29 help_understand 1.56 arrived_number 2.19 try_again 1.54
correct_solution 2.19 explain_steps 1.56 are_sure 2.19 thats _good 1.43
look_problem 2.18 took_arrive 1.56 sure_that 2.19 lets_break 1.37
great_effort 1.62 lets_step 1.51 correct_remember 1.38 glasses_water 1.3
lets_steps 1.55 better_understand 1.31 and_long 1.38 for_example 1.3
need_help 1.55 ones_place 1.31 digit_answer 1.38 times_equal 1.3
let_know 1.55 number_sides 1.31 answer_step 1.38 represents_glasses 1.29
@@ Acknowledgements
We’d like to thank Jiang Wu, Hannah Shuchhardt, and two anonymous individuals for their help and feedback on our work; the Stanford NLP group, Caleb Ziems, Joy He-Yueya, Gabriel Poesia, Myra Cheng, Kristina Gligorić, Ali Malik and Roma Patel for their feedback on the paper; Jesse Mu for pointing us to Sankey diagrams; and IMK for the Bridge inspiration.
@@ D.1 No Decision-Making Condition
No Decision-Making Prompt for gpt-4 and gpt-3.5-turbo
You are an experienced elementary math teacher and you are going to respond to a student’s mistake in a useful and caring way. The problem your student is solving is on topic: {lesson_topic}. {c_h} tutor (maximum one sentence):
@@
No Decision-Making Prompt for llama-2
### System: You are an experienced elementary math teacher and you are going to respond to a student’s mistake in a useful and caring way. ### User: Lesson topic: {lesson_topic} Conversation: {c_h} ### Assistant: tutor (maximum one sentence):
@@ D.2 Expert Decision-Making Condition
Decision-Making Prompt for gpt-4 and gpt-3.5-turbo
You are an experienced elementary math teacher and you are going to respond to a student’s mistake in a useful and caring way. The problem your student is solving is on topic: {lesson_topic}. {e} {z_what} in order to {z_why}. {c_h} tutor (maximum one sentence):
@@
Decision-Making Prompt for llama-2
### System: You are an experienced elementary math teacher and you are going to respond to a student’s mistake in a useful and caring way. ### User: {e} {z_what} in order to {z_why}. Lesson topic: {lesson_topic} Conversation: {c_h} ### Assistant: tutor (maximum one sentence):
@@ D.3 Self decision-making condition
Determine Error ($e$) with gpt-4 and gpt-3.5-turbo.
You are an experienced elementary math teacher. Your task is to read a conversation snippet of a tutoring session between a student and tutor, and determine what type of error the student makes in the conversation. We have a list of common errors that students make in math, which you can pick from. We also give you the option to write in your own error type if none of the options apply. Error list: 0. Student does not seem to understand or guessed the answer. 1. Student misinterpreted the question. 2. Student made a careless mistake. 3. Student has the right idea, but is not quite there. 4. Student’s answer is not precise enough or the tutor is being too picky about the form of the student’s answer. 5. None of the above, but I have a different description (please specify in your reasoning). 6. Not sure, but I’m going to try to diagnose the student. Here is the conversation snippet: Lesson topic: {lesson_topic} Conversation: {c_h} Why do you think the student made this mistake? Pick an option number from the error list and provide the reason behind your choice. Format your answer as: [{"answer": #, "reason": "write out your reason for picking # here"}]
@@
Determine Error ($e$) with llama-2.
### System: You are an experienced elementary math teacher. Your task is to read a conversation snippet of a tutoring session between a student and tutor, and determine what type of error the student makes in the conversation. We have a list of common errors that students make in math, which you can pick from. We also give you the option to write in your own error type if none of the options apply. Error list: 0. Student does not seem to understand or guessed the answer. 1. Student misinterpreted the question. 2. Student made a careless mistake. 3. Student has the right idea, but is not quite there. 4. Student’s answer is not precise enough or the tutor is being too picky about the form of the student’s answer. 5. None of the above, but I have a different description (please specify in your reasoning). 6. Not sure, but I’m going to try to diagnose the student. Format your answer as: [{"answer": #, "reason": "write out your reason for picking # here"}] ### User: Lesson topic: {lesson_topic} Conversation: {c_h} ### Assistant: [{"answer":
@@
Determine Strategy and Intention ($z_{\text{what}},z_{\text{why}}$) with gpt-4 and gpt-3.5-turbo.
You are an experienced elementary math teacher. Your task is to read a conversation snippet of a tutoring session between a student and tutor where a student makes a mistake. You should then determine what strategy you want to use to remediate the student’s error, and state your intention in using that strategy. We have a list of common strategies and intentions that teachers use, which you can pick from. We also give you the option to write in your own strategy or intention if none of the options apply. Strategies: 0. Explain a concept 1. Ask a question 2. Provide a hint 3. Provide a strategy 4. Provide a worked example 5. Provide a minor correction 6. Provide a similar problem 7. Simplify the question 8. Affirm the correct answer 9. Encourage the student 10. Other (please specify in your reasoning) Intentions: 0. Motivate the student 1. Get the student to elaborate their answer 2. Correct the student’s mistake 3. Hint at the student’s mistake 4. Clarify a student’s misunderstanding 5. Help the student understand the lesson topic or solution strategy 6. Diagnose the student’s mistake 7. Support the student in their thinking or problem-solving 8. Explain the student’s mistake (eg. what is wrong in their answer or why is it incorrect) 9. Signal to the student that they have solved or not solved the problem 10. Other (please specify in your reasoning) Here is the conversation snippet: Lesson topic: {lesson_topic} Conversation: {c_h} How would you remediate the student’s error and why? Pick the option number from the list of strategies and intentions and provide the reason behind your choices. Format your answer as: [{"strategy": #, "intention": #, "reason": "write out your reason for picking that strategy and intention"}]
@@
Determine Strategy and Intention ($z_{\text{what}},z_{\text{why}}$) with llama-2.
### System: You are an experienced elementary math teacher. Your task is to read a conversation snippet of a tutoring session between a student and tutor where a student makes a mistake. You should then determine what strategy you want to use to remediate the student’s error, and state your intention in using that strategy. We have a list of common strategies and intentions that teachers use, which you can pick from. We also give you the option to write in your own strategy or intention if none of the options apply. Strategies: 0. Explain a concept 1. Ask a question 2. Provide a hint 3. Provide a strategy 4. Provide a worked example 5. Provide a minor correction 6. Provide a similar problem 7. Simplify the question 8. Affirm the correct answer 9. Encourage the student 10. Other (please specify in your reasoning) Intentions: 0. Motivate the student 1. Get the student to elaborate their answer 2. Correct the student’s mistake 3. Hint at the student’s mistake 4. Clarify a student’s misunderstanding 5. Help the student understand the lesson topic or solution strategy 6. Diagnose the student’s mistake 7. Support the student in their thinking or problem-solving 8. Explain the student’s mistake (eg. what is wrong in their answer or why is it incorrect) 9. Signal to the student that they have solved or not solved the problem 10. Other (please specify in your reasoning) Format your answer as: [{"strategy": #, "intention": #, "reason": "write out your reason for picking # here"}] ### User: Lesson topic: {lesson_topic} Conversation: {c_h} ### Assistant: [{"strategy":
@@ Appendix F Lexical analysis
effort_remember 2.34 can_explain 2.14 actually_correct 2.9 thats_close 1.85
lets_focus 1.32 great_start 1.94 correct_answer 1.96 example_help 1.69
carry_tens 1.31 can_tell 1.88 job_attempting 1.42 can_think 1.51
focus_question 1.31 explain_got 1.53 small_mistake 1.42 can_try 1.51
clarify_mean 1.31 got_answer 1.42 attempting_problem 1.42 glasses_water 1.47
@@
user_lesson 8.02 lets_closer 3.73 user_student 5.79 lets_closer 3.27
user_tutor 4.27 closer_look 3.73 student_responds 4.65 closer_look 3.27
respond_students 3.62 look_problem 2.51 student_understand 4.0 right_track 2.49
mistake_useful 3.54 look_expression 1.56 response_provide 3.85 youre_right 2.09
going_respond 3.52 groups_objects 1.56 provide_hint 3.09 look_answer 1.7